OpenStudy (anonymous):

How do you implicitly differentiate (sinx+cosy)^2=1?

OpenStudy (zarkon):

where are you stuck?

OpenStudy (anonymous):

I tried using the chain rule for it, but my answer doesnt simplify at all :s

OpenStudy (zarkon):

why don't you show us what you tried.

OpenStudy (anonymous):

2(sinx+cosy)*(cosx-siny y') = 0 I multiplied them together and got y' = -2cosxcosy - 2sinxcosx/ -2(sinxsiny - cosysiny) so from there im stuck

OpenStudy (zarkon):

\[2(\sin(x)+\cos(y))*(\cos(x)-\sin(y) y') = 0 \] \[(\sin(x)+\cos(y))\cos(x)-(\sin(x)+\cos(y))\sin(y) y' = 0 \] \[(\sin(x)+\cos(y))\cos(x)=(\sin(x)+\cos(y))\sin(y) y' \] \[y'=\frac{(\sin(x)+\cos(y))\cos(x)}{(\sin(x)+\cos(y))\sin(y)} \] \[y'=\frac{\cos(x)}{\sin(y)}\]

OpenStudy (turingtest):

he divided by 2 (which drops out because t's all =0) then distributed sinx+cosy to cosx and -sinx in the next set of parentheses

OpenStudy (turingtest):

he divided by 2 (which drops out because t's all =0) then distributed (sinx+cosy) to cosx and -sinx in the next set of parentheses

OpenStudy (anonymous):

okay i kinda get that, but when you distribute the (sinx+cosy) to cosx and -sinx to the next would you have to foil them out? i end up getting sinxcosx -sinysinx + cosxcosy -siny y' cosy = 0

OpenStudy (turingtest):

to make it more clear: let\[\sin x+\cos y=a\]we then have\[a(\cos x-\sin yy')\]distributing we get\[a\cos x-a\sin yy'\]subbing back in \(\sin x+\cos y=a\) we get Zarkon's answer

OpenStudy (anonymous):

oh okay, so he didnt completely distribute them out but kept in a(cosx - sinyy') form i think i get it now, thanks !!

OpenStudy (turingtest):

Yeah, Zarkon is not so much about explanation, but he will answer the most difficult question with ease. He leaves the groundlings like me to explain his techniques.

OpenStudy (anonymous):

haha yeah, thanks for the explanation though because i find that more important than just the answer!

OpenStudy (turingtest):

that's my philosophy as well, so I'm happy to oblige :)